Summary of First Lecture on Dasein

Greg Hudson

November 18, 1994

Dasein is the word we use to talk about people, in order to avoid tangles with the English language. In order to take a realistic look at Dasein, we must reveal the identity of Dsein through Dasein’s engagement in projects and Dasein’s role in those projects. In fact, since the projects exist independent of Dasein, we say that Dasein is “thrown” into various projects, and “thrownness” replaces “involvement.” (This method of observation is interesting in that it resembles behaviorism at a first glance, but the principles of observation are much less mechanistic and reductionist.)

As an exmaple, Consider the project of learning French. The reductionist description is that we memorize a long list of words, learn rules of grammar, and thus learn how to make sentences. Actually, however, we learn to internlize knowledge of French when we gain familiarity, so that we can forget about the details such as rules of grammar. In a way, we learn to forget. In this way, learning structures itself into a repetition of the factical component of a phenomenon, which we call Wiederholung, and, following Wiederholung, an authentic familiarity with what was repeated. It is interesting that people often confuse spontaneity with innocence; spontaneity actually comes with great familiarity with a phenomenon.

The kinds of ontological phonemona we derive while studying the phenomenology of Dasein are called existentials. Some existentials are familiarity and disclosure. Disclosure has ontic modes of understanding and of emotional disclosure, which are equiprimordial and both important in becoming familiar with a situation.

This approach to looking at learning opposes a couple of common myths: the myth of rules, which states that we learn by learning rules and applying those rules deductively to various situations; and the myth of inference frmo general to particular, which states that we commonly learn by collecting data and verifying hypotheses about the data. Ordinarily, learning is not like this.